simplify the expression

cos(p+z)-cos(p-z)
Answer #1

Formula Used: (i) cos (P+z) i) cos (P-2) = cos(P) COS (2.) - Sin (P) sin(2) = cos(P) Cos(2) + sin(P) sin(2) Answer: COSCP+2)